Extend the transformation matrices in this lecture (scale, shear, and rotation) to 3 x 3 transformation matrices.8 Extend the transformation matrices in this lecture (scale, shear, and rotation) to 3 × 3 transformation matrices.
Answer #1

Date Page No. Extend transfarma:tion Lecture(scale, shear, and rotation). the matrices in. this 3x3 transfor mation to matric

Date Page No scalling_ transormation mat transformaHon matrices- extend. the to 3.x3 X Sx Y Sx x Sy C 1 The transformahon ob

Datend. Page No Shearing X - O y x = x + 5hy.Y (,2) (3,2) (2, o) (0,0) (0,0) (ariginal abject). (object afterx Shear) X 1 Sh

Page No is done it such Rotation Retation Gü. as angle following directions- an the help. the in clockwise andi - clockwise/c

Date Page No. A X = x Cos COs H y = ySin » Sin d= x xz x Cos (+9) Cos(a+O) = _A H X Sin (p+) = D. - H equation. from fCos(A+Page fromeq and be Linal equations will XSin + Y Cos O. Sin G ) Cos & Cos & Sin & into. extended transformation matices. whic